OpenStudy (anonymous):

Consider the parabola y = 4x − x2. (a) Find the slope of the tangent line to the parabola at the point (1, 3).

OpenStudy (helder_edwin):

do u know derivatives?

OpenStudy (anonymous):

i'm just getting into them

OpenStudy (anonymous):

i know what they are

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

just plug in 1 and get 2

OpenStudy (anonymous):

(b) Find an equation of the tangent line in part (a).

OpenStudy (helder_edwin):

ok u have \[ \large f(x)=4x-x^2 \] find \(f'(x)\)

OpenStudy (helder_edwin):

then the tangent line at (1,3) will be \[ \large y-3=f'(1)\cdot(x-1) \]

OpenStudy (anonymous):

how did you get there?

OpenStudy (helder_edwin):

to define a line u need either two points or one point and the slope. in this case u have the point (1,3) and the slope of the tangent which is the derivative.

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

did u get \(f'(1)=\)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it 0?

OpenStudy (helder_edwin):

\[ \large f(x)=4x-x^2 \] then \[ \large f'(x)=4-2x \] so \[ \large f'(1)=4-2\cdot1=2 \]

OpenStudy (anonymous):

ohh

OpenStudy (helder_edwin):

this is the slope of the tangent line

OpenStudy (anonymous):

ok